# supplementary

### v3

@@ 4.1 Emulating AI representatives
Ballot aggregation. For the US elections, majority determines the winner. For the two voting scenarios of participatory budgeting, the generated ballots by the AI representatives are aggregated using either the utilitarian greedy method or the method of equal shares. Utilitarian greedy simply selects the next most popular project, the one with the most received votes, as long as the available budget is not exhausted. Equal shares provides a proportional representation of voters’ preferences by dividing the budget equally among voters as endowments. This part of the budget can only be used to fund projects that the voter has voted for. The method goes through all project options, beginning from the projects with the highest number of votes. It selects a project if it can be funded using the budget shares of those who voted for the project. A full explanation of the is out of the scope of this article and can be found in earlier work [38, 48, 49]. In practice, the method of equal shares is likely to sacrifice an expensive popular project to select instead several low-cost projects, which, in combination, satisfy the preferences of more voters [39, 22]. As a result, equal shares tends to elect more winning projects. Because of this effect, it likely that consistency measurements based on the Jaccard similarity yield higher values for equal shares. This is the reason we control for the number of winning projects in equal shares by counting a subset of the most popular winning projects which is equal in number with the winners of the utilitarian greedy method.
Emulated voting processes with human voters and AI representatives. We emulate the following voting processes: (i) Human voters exclusively with 100% turnout. (ii) Human voters exclusively with varying turnout in the range [10%,100%] with a step of 10%. (iii) Mixed populations of human voters and AI representatives assessed of whether they can recover the consistency of voting outcomes lost by different human voters turnouts in the range [10%,100%] with a step of 10%. The ratio of AI representatives for the abstained voters varies in the range [25%,100%] with a step of 25%. (iv) AI representatives exclusively.
Each scenario of voters turnout is repeated 10 times by selecting a random subset of voters who abstain. For each turnout level, the AI representation of abstained voters is also repeated 10 times by selecting a random subset of the abstained voters to represent by AI.
Data collection infrastructure. Generative AI choices are collected via API prompts to large language models from 16th of June 2023 to 8th of November 2023. We prompted the large language models using the zero-shot learning feature [50], which does not require any specific fine-tuning. We use chain of thought prompting [55] to provide a comprehensive and systematic flow of information for better interpretability. The large language models run with temperature settings of 0.2 to 0 (20 runs for each temperature setting), for a focused decision with less randomness [56]. As we are dealing with a significantly large decision space, particularly the 33 projects in the actual voting, running with a high-temperature setting can lead to random decision-making. To maintain consistency, we retain the projects that are approved at the deterministic temperature setting (the responses are easily reproducible at temperature setting = 0). Furthermore, if a project is selected both in 0.1 and 0.2, we include the approval for that project in the ballot to also incorporate diversity. In the case of score voting, we take the frequency of the distinct scores assigned to a project across different temperature setting and add the score with the highest frequency. However, for the actual City Idea scenario for which we emulate cumulative voting, the total score provided in a ballot should be lower than the total number of points. Therefore, it would not be be accurate to derive a score for a project from multiple outputs corresponding to different temperature settings, hence we conducted experiments only with the temperature setting of 0.2. The decoding algorithm for the large language model is set to the nucleus sampling approach [55]. Compared to purely random sampling, it avoids generating outliers or improbable sequences of choices.
For the predictive AI choices, we use neural networks with 3 hidden layers and the relu [57] activation function for the sub-networks in the neural network. We employ the ensemble method [58] of designing multiple neural networks with different parameters and merging in the final layer with cross entropy loss [59]. This helps to generalize on different sizes of datasets and improve the prediction scores on unseen data.
@@ 4.2 Voting datasets
Table 1 outlines the settings of the emulated voting scenarios.
US elections. The 2012, 2016, and 2020 survey waves of the American National Election Study (ANES) [60] are used. They encompass 4,500 votes together with the respective voter information for each of these three years. This mainly includes socio-demographic characteristics: (i) racial/ethnic self-identification, (ii) gender, (iii) age, (iv) conservative-liberal ideological self-placement, (v) party identification (political belief), (vi) political interest, (vii) church attendance, (viii) discussion of politics with family/friends, (ix) feelings of patriotism associated with the American flag, and (x) state of residence. A total of 9 elections are emulated using Llama, GPT-3.5 and the predictive ML model based on single-choice ballots, with winners decided by majority for the three years. We extracted information from 20,650 voters of the ANES dataset [60] and after pre-processing (removing voters with missing values in more than 5 of the personal traits), a set of 17,010 voters is used. The examples of the prompts used to generate the AI choices can be found in Table LABEL:table:Anes.
City Idea: Participatory budgeting campaign in Aarau, Switzerland. The data from a recent innovative participatory budgeting field study are used [32, 39], which was conducted with ethical approval from University of Fribourg (#2021-680). It run in 2023 and is rigorously designed to assess the application of equal shares for the first time in real world, in combination with cumulative voting, using the open-source Stanford Participatory Budgeting platform [33]. As such, the field study includes a survey conducted before voting linking the choices of survey respondents and voters. We use the following personal human traits from the survey (more information in Tables LABEL:table:prologue1 and LABEL:table:prologue2): (i) 9 key socio-demographic characteristics (e.g., age, citizenship, education) and 2 political interests (political beliefs and trust in democracy), (ii) preferences for 9 different types of projects and 6 beneficiaries and (iii) 4 types of preferences for qualities of the voting outcome.
Table 1: The studied dimensions across three real-world voting scenarios. They provide the necessary diversity to generalize the findings of this study as they include different scales in number of voters, different ballot formats and aggregation methods, low and high numbers of alternatives, different personal human traits for studying a broad spectrum of biases including both generative and predictive AI methods.
@@
Socio-demographics
Preference for voting outcome qualities
Project preferences
Political interests
@@
Voters $\sim$17,000 (across 3 years) 3,134 505
Emulated elections 9 91 27
@@
    Mock Voting: Five hypothetical costed projects belonging in different categories are put for mock voting as part of the initial survey. Table LABEL:table:prologueS illustrates the project alternatives and their cost. The choice of 3134 voters over the same alternatives is tested with three different ballot formats in a sequence, starting with the simplest one of single choice to the most complex ones of approvals and score voting. This allows us to emulate a total of 45 elections = 3 ballot formats x 2 ballot aggregation methods x 3 AI models x 5 personal traits + 1 for the predictive ML model.
@@
    Actual Voting: Using the Stanford Participatory Budgeting platform [33], 1703 voters cast their vote using cumulative ballots by distributing 10 points to at least 3 projects of their preference, out of 33 projects in total (see Table LABEL:table:prologueA for project descriptions). A subset of 505 of these voters, which participated in the initial survey and provided their personal human traits, are used to construct the AI representatives. The ballot formats of single choice and approvals are derived from the cumulative ballots by taking the project with the most points and the projects that received any point respectively. This allows us to emulate a total of 91 elections = 3 ballot formats x 2 ballot aggregation methods x 3 AI models x 5 personal traits.
The examples of the prompts used to generate the AI choices in both mock and actual voting can be found in Table LABEL:table:prologuePersona.
@@ 4.3 Evaluation of choices by AI representatives
The emulated elections with AI representatives are compared to the real-world elections of human voters at two levels: (i) individual choice, i.e. the ballots, and (ii) collective choice, i.e. the resulting voting outcomes. Consistency is the main measure of assessment.
Consistency of individual choice. Consistency is assessed at two levels:
@@
    Human-AI consistency of individual choice: This is the Jaccard similarity [61] of individual choices made by AI representatives with the original choices of the humans they represent. For single-choice ballots, this is either 1 (consistent) or 0 (inconsistent) for each voter and the respective AI representative. For approval voting, the AI and human choices are binary sequences representing the ballot, containing 1 (approved) or 0 (not approved) for each alternative. The Jaccard similarity between these sequences is the ratio between the number of alternatives approved by both the human and the AI representative over the number of alternatives approved by at least one of them. For score and cumulative voting, projects are ranked based on the scores and points received respectively. Each top-k project of a human voter is compared to the respective top-k project by its AI representative. Matching is indicated by 1 (consistent), otherwise 0 (inconsistent) indicates no matching. The mean number of matchings among all ranked projects measures the consistency for these ballot formats.
@@
[…]
Consistency of collective choice. It is defined for two voting processes that elect winners from the same alternatives, but have different voters (human, AI representatives, mixed), and, as a result, can yield different voting outcomes. The voting outcomes with the winners are first calculated using one of the ballot aggregation methods, i.e. majority, utilitarian greedy or equal shares. These outcomes are represented as binary sequences containing either 0 (not winner) or 1 (winner) for each alternative. Consistency is calculated using the Jaccard similarity [61] between these binary sequences.
@@
Consistency recovery in collective choice with AI representatives. It is determined here for voting processes with varying voters turnout, in which abstained voters result in consistency loss, which can be recovered if a portion of these abstained voters are represented by AI. It is measured as follows:
$\frac{consistency\ [\text{human\ voters - abstained\ voters + AI\ representatives}]-consistency\ [\text{human\ voters - abstained\ voters}]}{1-consistency\ [\text{human\ voters - abstained\ voters}]},$
@@
where the voters turnout human voters - abstained votershuman voters + abstained $\frac{\text{human\ voters - abstained\ voters}}{\text{human\ voters + abstained\ voters}}$ varies in the range [10%,90%] with a step of 10%, and AI representation AI representativesabstained $\frac{\text{AI\ representatives}}{\text{abstained\ voters}}$ varies in the range [25%,100%] with a step of 25%.
4.4 Human cognitive biases in AI collective decision making
Human choices are significantly influenced by potential cognitive biases that are often a manifestation of socio-economic characteristics, conditions of life quality, satisfaction with the available public amenities and the overall life experiences of an individual [41]. Here we map self-reported personal human traits with which large language models are prompt to potential underlying human cognitive biases, as a way to explore whether they are reinforced by the language models and become more likely to manifest in the AI choices. Figure 1c outlines the mapping we study based on a systematic and comprehensive review of relevant literature. The following types of biases are determined:
Time-discounting biases. These are characterized by the tendency to receive immediate gratification over a larger but future reward. Projects related to the public space, culture, or welfare are often one-off projects, such as an annual festival, a cinema night, or educating asylum children (alternatives proposed for the participatory budgeting campaign in Aarau [32]), where expected benefits are experienced in the short term. Therefore, such project are subject of time-discounting biases [41]. In contrast, the construction of a road, for example, incurs investment costs several years before the actual project implementation. Such investments create a sustainable reform that benefits all in the long run. Even though 72% of the public transportation infrastructure in European cities ends up in cost overruns, people still vote for these projects due to an underlying necessary optimism to improve transport [41, 39]. We refer to this economic discounting over a long-term sustainability impact as ‘optimism’ [62, 63].
Surrogation biases. This reflects how humans favor simpler measures to assess the impact over ones that are more precise and harder to evaluate. Korteling et al. [41] argue that these biases manifest when deciding projects with large societal impact such as health or education, while their outcome is subject of different satisfaction levels among citizens. The project outcomes may be perceived as successful by part of society using easy-to-evaluate metrics instead of looking at long-term effects on the community. For instance, a timely vaccination drive may be preferred over significant changes to vaccination protocols or health insurance policies covering vaccination. Hence these projects are likely to be more preferred as they come with more intuitive ways to assess for the broader population. This is reflected by the average winning rate of the health and education-related projects of 38.1% and 36.2% respectively in Poland [39], where participatory campaigns have been actively hosted in the last decade.
Conformity biases. These biases arise out of group pressure under which people make decisions to be socially desirable [40, 42]. It is argued that a conformity bias may induce voting for green alternatives [41]. Green-themed participatory budgeting campaigns have been adopted in European cities such as Lisbon [64], to promote green initiatives, aligning to a culture for more sustainable behavior. Poland runs participatory budgeting campaigns at large scale, which include environment and urban greenery projects. These are within the top-5 most popular projects with an average of 22.5% and 26.5% respectively [39]. Even in Aarau we observe the same trend wherein, environmental friendly projects account for the top-10 most popular projects [32].
Affect heuristic biases. This defines the tendency to make decisions based on what intuitively or emotionally feels right. Affect biases have been studied to analyze the inclusive attitudes most people show towards elderly people [65]. Similar biases also manifest in welfare of children and in inter-generational communication [41, 66]. In Aarau, we observe that 71.3% of the voters prefer projects for younger and elderly people.
Biases for altruism and egotism. Individual interest is often in conflict with the community interest in participatory and collective decision-making processes. While voting, the intrinsic altruism of citizens strongly influences choices. As a result, altruism and egotism are influential for the fairness of voting outcomes and how these outcomes benefit the city in overall [67, 41, 66]. In Aarau, we observe that 67.1% of the voters, who prefer better representation in the outcome are prosocial and prioritize the city-wide benefit (altruism bias) over individual benefit (egotism bias).
Unconscious biases. Human choices are influenced by socio-economic and demographic traits such as race, ethnicity, citizenship, household size and income [68]. Specifically, political ideology and belief shape to a high degree decisions for candidates in elections [69].
@@ 4.5 Explainability of generative AI voting
The consistency of the individual AI choices (see Section 4.3) is modeled as the dependent variable in a predictive machine learning framework. Causal relationships that explain how personal human traits (independent variables) contribute to consistency are studied. We model the problem of explaining inconsistencies as a classification problem, where 10 uniform consistency levels are defined as $[0.0,0.1],(0.1,0.2],...,(0.9,1.0]$. Further details about how we account for imbalances of features, their collinearity and hyperparameter optimization of the model are illustrated in Section LABEL:sec:ml and Table LABEL:tab:set. We also assess prediction models with different dependent variables, see Figure LABEL:fig:explaination-top-choice for more details.
Explainability of choices. We introduce a two-dimensional feature importance analysis framework to determine the impact of the personal human traits on the consistency of individual choice. We use model agnostic Shapley Additive Explanations [70] to extract individual contributions of each trait. Results are shown in Figure 5. Feature ablation study [71] is used to calculate the error (loss) in the overall prediction accuracy of the model when a feature is removed (results in Table LABEL:table:error).

### v4

@@ 4.1 Emulating AI representatives
Ballot aggregation. For the US elections, majority determines the winner. For the two voting scenarios of participatory budgeting, the generated ballots by the AI representatives are aggregated using either the utilitarian greedy method [69] or the method of equal shares [39]. Utilitarian greedy simply selects the next most popular project, the one with the most received votes, as long as the available budget is not exhausted. Equal shares provides a proportional representation of voters’ preferences by dividing the budget equally among voters as endowments. This part of the budget can only be used to fund projects that the voter has voted for. The method goes through all project options, beginning from the projects with the highest number of votes. It selects a project if it can be funded using the budget shares of those who voted for the project. A full explanation is out of the scope of this article and can be found in earlier work [39, 57, 58]. In practice, the method of equal shares is likely to sacrifice an expensive popular project to select instead several low-cost projects, which, in combination, satisfy the preferences of more voters [27, 28]. As a result, equal shares tends to elect more winning projects. Because of this effect, it is likely that consistency measurements based on the pairwise similarity yield higher values for equal shares. This is the reason we control for the number of winning projects in equal shares by counting a subset of the most popular winning projects which is equal in number with the winners of the utilitarian greedy method. We also use the Phragmén’s proportional aggregation rule [40]. Phragmén’s method balances fairness and representation between groups, whereas the Method of Equal Shares emphasizes fair representation within groups, ensuring that at least one voter from each group is represented [70].
Modeling abstaining voters. We model three types of abstaining voters: (i) Those whose low digital skills may be a barrier to effectively use digital platforms to participate online and vote [12, 52], which often results in abstaining from participatory processes [8, 53, 54, 42]. (ii) Those with lower political engagement [41, 51] and (iii) distrust to institutions [55, 56, 42] often choose to abstain from voting. We identify pre- and post-voting survey questions participatory budgeting campaign of City Idea (see Table LABEL:table:prologue3LABEL:table:prologue4 and LABEL:table:prologue5) that are proxies to determine potential voters with one of the three abstaining profiles [54, 71]. We divide voters to two quartiles based on their responses to distinguish the voters that are likely to abstain and the ones that are likely to participate. This results in voters with low digital literacy, engagement and trust that characterize 48.25%, 36.12% and 27.36% of the population respectively.
Emulating different participation modalities. Four participation modalities are studied: (i) Human voters exclusively with 100% of voters’ turnout. (ii) Human voters exclusively with varying turnout levels in the range [10%,100%] with a step of 25%. The maximum number of abstaining voters is either the total voters (baseline in Figure 4b) or the number of voters with low digital literacy, engagement and trust as determined by the abstaining models. (iii) Mixed populations of human voters and AI representatives of abstaining voters in the range [25%,100%] with a step of 25%. (iv) AI representations exclusively.
A random control to rigorously assess AI representation of abstaining voters. The consistency loss by a number of abstaining voters belonging to one of the three groups of voters with low digital literacy, engagement and trust, as well as the recovery by their AI representation, is compared with the mean consistency loss and recovery of 30 groups of random voters, each of size equal to the abstaining voters (see Figure 4). For each turnout level of the abstaining voters, the process is also repeated 30 times by selecting a random subset of abstaining voters to represent by AI.
Data collection infrastructure. Generative AI choices are collected via API prompts to large language models from 16th of June 2023 to 8th of November 2023. We prompted the large language models using the zero-shot learning feature [59], which does not require any specific fine-tuning. We use chain of thought prompting [72] along with context-based prompting [73] to provide a comprehensive and systematic flow of information for better interpretability. A detailed explanation for the prompt designing is provided in Section LABEL:sec:prompt.
@@ 4.2 Voting datasets
Table 1 outlines the characteristics of the emulated voting scenarios.
US elections. The 2012, 2016, and 2020 survey waves of the American National Election Study (ANES) [74] are used. The dataset for three years contains 20,650 voters together with the respective voter socio-demographic characteristics for each of these three years: (i) racial/ethnic self-identification, (ii) gender, (iii) age, (iv) conservative-liberal ideological self-placement, (v) party identification (political belief), (vi) political interest, (vii) church attendance, (viii) discussion of politics with family/friends, (ix) feelings of patriotism associated with the American flag, and (x) state of residence. A total of 18 elections are simulated using two combinations of human traits and three AI models, including Llama, GPT-3.5, and a predictive ML model based on single-choice ballots, with winners determined by majority vote over three years. We extracted information from 20,650 voters of the ANES dataset [74] and after pre-processing (removing voters with missing values in more than 5 of the personal traits), a set of 17,010 voters is used. The examples of the prompts used to generate the AI choices can be found in Table LABEL:table:Anes.
City Idea: Participatory budgeting campaign in Aarau, Switzerland. The data from a recent innovative participatory budgeting campaign are used [33, 27], which was conducted with ethical approval from University of Fribourg (#2021-680). It run in 2023 and is rigorously designed to assess the application of equal shares for the first time in real world, in combination with cumulative voting, using the open-source Stanford Participatory Budgeting platform [34]. As such, the field study includes a survey conducted before voting linking the choices of survey respondents and voters. We use the following personal human traits from the survey (more information in Tables LABEL:table:prologue1 and LABEL:table:prologue2): (i) 9 key socio-demographic characteristics (e.g., age, citizenship, education) and 2 political interests (political beliefs and trust in democracy), (ii) preferences for 9 different types of projects and 6 beneficiaries and (iii) 4 types of preferences for qualities of the voting outcome. Sc
Table 1: The studied dimensions across three real-world voting scenarios. They provide the necessary diversity to generalize the findings of this study as they include different number of voters, different ballot formats and aggregation methods, low and high numbers of alternatives, different personal human traits for studying a broad spectrum of biases including both generative and predictive AI methods.
@@
Socio-demographics, Outcome Expectations, Project Preferences, Political Interests
Socio-demographics, Outcome Expectations, Project Preferences
Socio-demographics, Project Preferences, Political Interests
Socio-demographics, Outcome Expectations, Political Interests
Socio-demographics, Political Interests
Socio-demographics, Political Interests (only 1 feature)
@@
Voters $\sim$17,010 (across 3 years) 3,314 505
Emulated elections 18 180 108
@@
    Survey Voting: Five hypothetical costed projects belonging in different categories are put for choice as part of the initial survey. Table LABEL:table:prologueS illustrates the project alternatives and their cost. The choice of 3314 voters over the same alternatives is tested with three different ballot formats in a sequence, starting with the simplest one of single choice to the most complex ones of approvals and score voting. This allows us to emulate 180 elections = 3 ballot formats x 4 AI models x 5 combinations of personal traits x 3 ballot aggregation methods. An additional 9 elections were emulated using all human traits, the three ballot format - (single choice, approval and score) and greedy and equal shares aggregation for the GPT 4, only to evaluate the differences with GPT 3.5. There were no difference in collective consistency and mean of the individual consistencies varies with $\pm$0.05 when compared to that of GPT 3.5.
@@
    Actual Voting: Using the Stanford Participatory Budgeting platform [34], 1703 voters cast their vote using cumulative ballots by distributing 10 points to at least 3 projects of their preference, out of 33 projects in total (see Table LABEL:table:prologueA for project descriptions). A subset of 505 of these voters, which participated in the initial survey and provided their personal human traits, are used to construct the AI representatives. The ballot formats of single choice and approvals are derived from the cumulative ballots by taking the project with the most points and the projects that received any point respectively. This allows us to emulate 108 elections = 3 ballot formats x 4 AI models x 3 combinations of personal traits x 3 ballot aggregation methods.
The examples of the prompts used to generate the AI choices in both survey and actual voting can be found in Table LABEL:table:prologuePersona.
@@ 4.3 Evaluation of choices by AI representatives
The emulated elections with AI representatives are compared to the real-world elections of human voters at two levels: (i) individual choice, i.e. the ballots, and (ii) collective choice, i.e. the resulting voting outcomes. Consistency is the key assessment measure, derived from the accuracy of individual and collective AI choices compared to human decisions and the transitivity across different ballot formats (see Figure 1a).
Consistency of individual choice. Single-choice ballots for both AI and human voters are represented as binary sequences, where a value of 1 indicates approval of a specific project, and 0 denotes disapproval of all remaining alternatives. In approval voting, each alternative is assigned either 1 (approved) or 0 (not approved). In contrast, in score voting and cumulative voting each alternative receives a score or an number of distributed points (integer numbers) reflecting voter preference. To compare AI-generated and human choices, we employ a single method, the Condorcet pairwise comparison method [22, 13], which is a generic approach to characterize the overall similarity of two ballots (or voting outcomes). A preference matrix is constructed, where rows and columns correspond to alternatives, and each matrix element records the outcome of a pairwise comparison. If project $P_{i}$ is ranked higher than project $P_{j}$, or if $P_{i}$ is approved while $P_{j}$ is not, the corresponding matrix cell $P_{i}>P_{j}$ is assigned a value of 1; otherwise, it is set to 0. Ties are excluded from the analysis.
@@
    Human-AI consistency (accuracy) of individual choices: The human ballots serve as the reference point for evaluating the ones generated by the AI representatives, see Figure 1a. The elements of ‘1’ in the matrix of AI representatives that match the elements of ‘1’ in the matrix of human choices determine the consistency [13].
@@
    Consistency (transitivity) of AI and human individual choice across ballot formats: Ballot formats are standardized as follows: For cumulative/score vs. single-choice ballots, the highest-scoring projects are set to ‘1’ and the others to ‘0’. For cumulative/score vs. approval ballots, scored projects are set to ‘1’, while projects without score are set to ‘0’. The elements of ‘1’ and ‘0’ in the two matrices of the ballot formats that match determine the consistency.
We also compare the choices based on preference reordering using the Kemeny distance [75] as illustrated in Section LABEL:sec:supllemethod.
Consistency of collective choice. This follows the same approach of Condorcet pairwise comparisons for individual choices. However, before calculations of consistency are made, voting outcomes are turned into binary sequences to distinguish winners (‘1’) from losers (‘0’) as determined by a ballot aggregation method.
@@
Consistency recovery in collective choice with AI representatives. It is determined here for voting scenarios with varying voters turnout, in which abstained voters result in collective consistency loss, which can be recovered if a portion of these abstained voters are represented by AI. This recovery takes place at two levels: (i) False negative projects that are erroneously removed under abstaining but added back by AI representatives. (ii) False positive projects that are erroneously added under abstaining but correctly removed by AI representatives. Consistency recovery is measured as follows:
$\frac{consistency\ [\text{all\ human\ voters - abstained\ voters + AI\ representatives}]-consistency\ [\text{all\ human\ voters - abstained\ voters}]}{1-consistency\ [\text{all\ human\ voters - abstained\ voters}]},$
@@
where the voters turnout human voters - abstained votershuman voters + abstained $\frac{\text{human\ voters - abstained\ voters}}{\text{human\ voters + abstained\ voters}}$ varies in the range [20%,75%] with a step of 25%, and AI representation AI representativesabstained $\frac{\text{AI\ representatives}}{\text{abstained\ voters}}$ varies in the range [25%,100%] with a step of 25%.
@@ 4.4 Explainability of generative AI voting
The accuracy of the individual AI choices (see Section LABEL:sec:explain) with human choices as well as the transitivity of AI choices over different ballot formats are modeled as the dependent variable in a predictive machine learning framework. We study causal relationships explaining how personal human traits (independent variables) influence consistency (both accuracy and transitivity). We model the problem of explaining inconsistencies as a classification problem, where 10 uniform consistency levels are defined as the ranges $[0.0,0.1],(0.1,0.2],...,(0.9,1.0]$. Further details about how we account for imbalances of features, their colinearity and hyperparameter optimization of the model are illustrated in Section LABEL:sec:ml and Table LABEL:tab:set.
Explainability of choices. We introduce a two-dimensional feature importance analysis framework to determine the impact of the personal human traits on the consistency of individual choices. For a given performance of the prediction model, we employ explainable AI methods to analyze the contribution of each individual human trait (feature) to the outcome. The approach to enhance the performance (accuracy, precision, recall) of the prediction model is illustrated in Section LABEL:sec:ml. We then use the model agnostic Shapley Additive Explanations (SHAP) and Local Interpretable Model Agnostic Explanations (LIME) [43] to extract the individual contributions of each trait. Results are shown in Figure LABEL:fig:ML_suppleLABEL:fig:Lime_scoreLABEL:fig:Lime_approvalLABEL:fig:approval and Table LABEL:table:error. A feature ablation study [76] is used to calculate the error (loss) in the overall prediction accuracy of the model when a feature is removed (results in Table LABEL:table:error).